Determine the standard form of the equation of a hyperbola with vertices (2,0) and (-2,0) and passing through (4,3)
Answer by venugopalramana(3286)

Determine the standard form of the equation of a hyperbola with vertices (2,0) and (-2,0) and passing through (4,3)
THE EQN. OF A HYPERBOLA IN STD.FORM IS
(X-H)^2/A^2 - (Y-K)^2/B^2=1….
WHERE
(H,K) IS CENTRE…HERE WE HAVE CENTRE IS {(2-2)/2,(0+0)/2}=(0,0)..H=K=0
VERTICES ARE (-A,0),(A,0)...A=2
HENCE EQN.IS
X^2/2^2-Y^2/B^2=1
IT IS PASSING THROUGH (4,3)....HENCE
4^2/4-3^2/B^2=1
3=9/B^2
B^2=9/3=3
HENCE EQN.IS
X^2/4-Y^2/3=1